# Massless higher-spin supermultiplets in 5D harmonic superspace

### v1

@@ 1 Introduction
In the literature, there are several gauge-invariant formulations for massive higher-spin fields, see Pashnev:1989 ; Klishevich:1997pd ; Klishevich:1998yt ; Zinoviev:2001 ; Buchbinder:2006nu ; Metsaev ; Buchbinder:2008ss ; Lindwasser for an incomplete list of publications. Using such formulations, one may obtain on-shell supersymmetric massive higher-spin models by combining bosonic and fermionic models, as was achieved by Zinoviev and collaborators Zinoviev:2007js ; Buchbinder:2015mta ; Buchbinder:2017izy ; Buchbinder:2019dof . This approach is not efficient, however, for constructing off-shell gauge-invariant actions for massive supermultiplets. As is well known, off-shell supersymmetry has been crucial for deriving the most general supergravity-matter systems with four supercharges in $d=4$ and eight supercharges in $d=4,5,6$ dimensions, see WB ; GIOS ; FVP ; Lauria:2020rhc for comprehensive reviews. It is worth expecting that off-shell gauge-invariant formulations for massive supermultiplets may facilitate the problem of constructing consistent interactions for such fields. Is there a systematic procedure to derive such formulations?
@@
Off-shell massless higher-spin supermultiplets were constructed in four and three dimensions. In the four-dimensional case, the ${\cal N}=1$ supermultiplets were first derived in Minkowski superspace KSP ; KS (see BK for a pedagogical review) and then in anti-de Sitter (AdS) superspace KuzenkoSibiryakov94 ; Buchbinder:2018nkp . ${\cal N}=1$ supersymmetric models proposed in KuzenkoSibiryakov94 were subsequently used to derive off-shell ${\cal N}=2$ higher-spin supermultiplets in AdS4 Gates:1996my . [^2] Building on the models introduced in KuzenkoSibiryakov94 , the universal linearized action for massless gauge multiples of arbitrary superspin in AdS superspace was proposed in Ref. Gates:1996xs as a gauge theory of unconstrained superfields taking their values in the commutative algebra of analytic functions over a one-sheeted hyperboloid in ${\mathbb{R}}^{3,1}$. The action is invariant under ${\cal N}=2$ supersymmetry transformations which form a closed algebra off the mass shell. The higher-spin models introduced in Ref. Gates:1996my are invariant under ${\cal N}=2$ supersymmetry transformations which form a closed algebra in AdS4, but this algebra becomes open in a flat limit. [^3] The higher-spin ${\cal N}=2$ supermultiplets of Gates:1996my are obtained by combining two massless ${\cal N}=1$ supersymmetric models corresponding to superspins $s$ and $s+1/2$, with $s$ a positive integer, and thus they cannot be considered as genuine ${\cal N}=2$ off-shell supermultiplets. The closure of the algebra of ${\cal N}=2$ supersymmetry transformations in these models is a specific feature of the AdS4 supersymmetry, and it is also manifested in the case of ${\cal N}=2$ supersymmetric nonlinear sigma models in AdS4 realized in terms of ${\cal N}=1$ chiral superfields Butter:2011zt ; Butter:2011kf . Genuine off-shell massless ${\cal N}=2$ supermultiplets were constructed in Buchbinder:2021ite within the harmonic-superspace approach.
@@
In three dimensions, massless higher-spin supersymmetric models do not have any propagating degrees of freedom. However, these models can be combined with linearized superconformal higher-spin actions to generate off-shell topologically massive higher-spin supersymmetric theories, following the philosophy of topologically massive gauge theories WS ; JS ; DJT1 ; DJT2 ; DeserKay . Recently, there has been much progress in the construction of such off-shell theories with various supersymmetries. It is based on the approaches developed in Kuzenko:2016qwo for ${{\cal N}}=1$ and in KO for ${{\cal N}}=2$ Poincaré supersymmetries. They were later extended to $\rm AdS_{3}$ for the following cases: ${\cal N}=1$ Kuzenko:2018lru ; HK19 ; Kuzenko:2021hyd , ${\cal N}=(1,1)$ HKO and ${\cal N}=(2,0)$ HK18 . Linearized actions for $\cal N$-extended higher-spin superconformal gravity were derived in Buchbinder:2019yhl . [^4] In four dimensions, linearized superconformal higher-spin actions were derived in KMT ; KP ; Kuzenko:2020jie for ${\cal N}=1$ and Kuzenko:2021pqm for ${\cal N}>1$, including the discovery of ${\cal N}=2$ superconformal gravitino multiplet HKR . The ${\cal N}=2$ results of Kuzenko:2021pqm ; HKR were recast in terms of projective Kuzenko:2024vms and harmonic BIZ24 ; Ivanov:2024bsb prepotentials.
@@
The rest of the paper is organized as follows. In Sec. 2 we give a short overview of existing superfield approaches to 5D supergravity and discuss some key results which represent a background for further sections. In Sec. 3, we introduce prepotentials for 5D Poincaré supergravity in the harmonic superspace framework. We use these prepotentials for construction of the superfield action for linearized supergravity in flat Minkowski superspace and consider minimal hypermultiplet coupling to this supergravity background. This allows us to derive conserved current superfields for the $q$-hypermultiplet model. This section serves as a starting point for introduction of superfield prepotentials with the corresponding gauge transformations for higher-spin supermultiplets which are given in Sec. 4. There we construct gauge-invariant quadratic action for higher-spin supermultiplets in harmonic superspace and derive the corresponding superfield equations of motion. Next, we build a cubic interaction vertex for these supermultiplets with $q$-hypermultiplet. This vertex allows us to derive conserved higher currents in the $q$-hypermultiplet model. Section 5 is devoted to a discussion of other interesting open problems which may be addressed building upon the result of the present paper. In Appendix A we give a summary of our superspace notation and conventions, while Appendix B contains technical details of the derivation of component structure of higher-spin prepotentials in the Wess-Zumino gauge. Finally, in Appendix C, we present prepotentials and gauge transformations for 5D conformal supergravity.
@@ 2 Comments on 5D superfield supergravity
General ${\cal N}=1$ supergravity-matter systems in five dimensions are naturally described using the superconformal tensor calculus Ohashi1 ; Ohashi2 ; Ohashi3 ; Ohashi4 ; Bergshoeff1 ; Bergshoeff2 ; Bergshoeff3 , see also Lauria:2020rhc for a pedagogical review, [^6] In the literature, ${\cal N}=1$ supersymmetry in five dimensions is often denoted ${\cal N}=2$, see e.g. Lauria:2020rhc ; Ohashi1 ; Ohashi2 ; Ohashi3 ; Ohashi4 ; Bergshoeff1 ; Bergshoeff2 ; Bergshoeff3 , by analogy with the 4D ${\cal N}=2$ case to indicate that the number of supercharges is equal to eight. in which hypermultiplets are either on-shell or involve a gauged central charge. These hypermultiplet realizations cannot be used to provide an off-shell formulation for the most general locally supersymmetric sigma model. Such a sigma model formulation is known to require one to use off-shell hypermultiplets possessing an infinite number of auxiliary fields, which make them extremely difficult to work with at the component level. This technical problem was solved within the superspace approach to 5D ${\cal N}=1$ supergravity-matter systems KT-M_5D2 ; KT-M_5D3 ; KT-M08 by putting forward the concept of covariant projective multiplets. These supermultiplets are a curved-superspace extension of the 4D ${\cal N}=2$ and 5D ${\cal N}=1$ superconformal projective multiplets Kuzenko2006 ; K07 . The latter reduce to the off-shell projective multiplets pioneered by Lindström and Roček KLR ; LR1 ; LR2 in the 4D ${\cal N}=2$ super-Poincaré case and generalized to the cases of 5D ${\cal N}=1$ Poincaré and AdS supersymmetries in KL and Kuzenko:2007aj , respectively.
The powerful virtues of the superconformal tensor calculus and of the superfield supergravity are naturally combined in the framework of 5D ${\cal N}=1$ conformal superspace Butter:2014xxa which was obtained by gauging the 5D superconformal algebra in superspace. This work reproduced practically all off-shell constructions derived at the time and provided a plethora of new results. In particular, it provided covariant superspace realizations for all off-shell formulations of 5D ${\cal N}=1$ supergravity with the Weyl supermultiplet coupled to two compensating supermultiplets. [^7] Numerous applications of the 5D conformal superspace approach have appeared Hutomo:2022hdi ; Gold:2023dfe ; Gold:2023ymc ; Gold:2023ykx ; Gold:2025ttt . The supergravity equations of motion were obtained in Butter:2014xxa making use of the fact that the 5D ${\cal N}=1$ Weyl multiplet is described in superspace by a single unconstrained real prepotential, $\mathfrak{U}$, similarly to the 4D ${\cal N}=2$ case HST ; KT ; Butter:2010sc . Given a supersymmetric theory of (matter) superfields $\varphi^{i}$ coupled to the Weyl multiplet, the supercurrent of this theory is a (dimension-3 primary) real scalar superfield defined by
@@
In the case of 4D ${\cal N}=2$ supergravity, the unconstrained prepotential $\mathfrak{U}$ for the Weyl multiplet originates most elegantly KT ; Butter:2010sc within the harmonic superspace approach GIKOS ; Galperin:1987em ; Galperin:1987ek . The same approach has recently been used to construct linearized off-shell actions for 4D ${\cal N}=2$ massless higher-spin supermultiplets Buchbinder:2021ite ; Buchbinder:2022kzl ; Buchbinder:2022vra since the supergravity prepotentials in the harmonic superspace possess a natural generalization to higher-spin supermultiplets. Therefore, in the next section, we develop prepotentials for linearized ${\cal N}=1$ Poincaré supergravity in 5D harmonic superspace and construct corresponding quadratic action for the spin-2 supermultiplet. [^8] In this paper, a supermultiplet which contains bosonic fields of integer spins $s$ and s1s-1, as well as one Dirac-type fermionic field of spin s1/2s-1/2 on shell is referred to as a spin-$s$ supermultiplet. In Sec. 4, this formulation will be extended for description of higher-spin multiples in 5D harmonic superspace. The hypermultiplet coupling to the 5D Weyl multiplets is discussed in Appendix C.
@@
As mentioned above, off-shell formulations for 5D minimal supergravity are obtained by coupling the Weyl multiplet to two compensating supermultiplets. One of the compensators is a vector supermultiplet. In the harmonic superspace framework, prepotentials for linearized Weyl multiplet are given by the superfields $h^{++m}$, $h^{++\hat{\alpha}+}$ and $h^{(+4)}$ with constraints and gauge transformations described in Appendix C. The coupling to the vector multiplet compensator is accounted for by adding the prepotential $h^{++5}$ such that the covariant harmonic derivative becomes
@@
where $D^{++}$ and $D^{--}$ are flat harmonic derivatives and $\partial_{5}$ is the central charge operator. Functional form of $h^{++5}$ is fixed by the requirement that the operator (2) preserves the analyticity of a massive $q$-hypermultiplet (with a non-vanishing central charge).
@@
which are superconformal provided that $L_{\mathbf{a}}^{i}$ is primary and dimensionless. The action for 5D ${\cal N}=1$ Poincaré supergravity in this setting is proportional to the vector multiplet action with the wrong sign. Several equivalent expressions for the supergrvaity action exist in the literature, including the following: Butter:2014xxa .
@@
where $V_{ij}$ is Mezincescu’s prepotential for the vector multiplet, $W$ is the field strength of the vector multiplet, and $\nabla^{ij}=\nabla^{\hat{\alpha}(i}\nabla_{\hat{\alpha}}^{j)}$, see Butter:2014xxa for the technical details. The nonlinear multiplet appears in the action (4) implicitly in the sense that the constraints (3) express some component fields of the Weyl multiplet in terms of those belonging to the nonlinear multiplet.
In the harmonic superspace framework GIOS ; Galperin:1987ek , it is described by an analytic superfield $N^{++}$ satisfying the nonlinear constraint [^10] The relationship between $N^{++}$ and $L_{\mathbf{a}}^{i}$ is described in Butter:2015nza .

### v2

@@ 1 Introduction
In the literature, there are several gauge-invariant formulations for massive higher-spin fields, see Pashnev:1989 ; Klishevich:1997pd ; Klishevich:1998yt ; Zinoviev:2001 ; Buchbinder:2006nu ; Metsaev ; Buchbinder:2008ss ; Lindwasser for an incomplete list of publications. Using such formulations, one may obtain on-shell supersymmetric massive higher-spin models by combining bosonic and fermionic models, as was achieved by Zinoviev and collaborators Zinoviev:2007js ; Buchbinder:2015mta ; Buchbinder:2017izy ; Buchbinder:2019dof . This approach is not efficient, however, for constructing off-shell gauge-invariant actions for massive supermultiplets. As is well known, off-shell supersymmetry has been crucial for deriving the most general supergravity-matter systems with four supercharges in $d=4$ and eight supercharges in $d=4,5,6$ dimensions, see WB ; GIOS ; FVP ; Lauria:2020rhc for comprehensive reviews. Therefore one may expect that off-shell gauge-invariant formulations for massive supermultiplets may facilitate the problem of constructing consistent interactions for such fields. Is there a systematic procedure to derive such formulations?
@@
Off-shell massless higher-spin supermultiplets were constructed in four and three dimensions. In the four-dimensional case, the ${\cal N}=1$ supermultiplets were first derived in Minkowski superspace KSP ; KS (see BK for a pedagogical review) and then in anti-de Sitter (AdS) superspace KuzenkoSibiryakov94 ; Buchbinder:2018nkp . The ${\cal N}=1$ supersymmetric models proposed in KuzenkoSibiryakov94 were subsequently used to derive off-shell ${\cal N}=2$ higher-spin supermultiplets in AdS4 Gates:1996my . [^2] Building on the models introduced in KuzenkoSibiryakov94 , the universal linearized action for massless gauge multiplets of arbitrary superspin in AdS superspace was proposed in Ref. Gates:1996xs as a gauge theory of unconstrained superfields taking their values in the commutative algebra of analytic functions over a one-sheeted hyperboloid in ${\mathbb{R}}^{3,1}$. The action is invariant under ${\cal N}=2$ supersymmetry transformations which form a closed algebra off the mass shell. The higher-spin models introduced in Ref. Gates:1996my are invariant under ${\cal N}=2$ supersymmetry transformations which form a closed algebra in AdS4, but this algebra becomes open in a flat limit. [^3] The higher-spin ${\cal N}=2$ supermultiplets of Gates:1996my are obtained by combining two massless ${\cal N}=1$ supersymmetric models corresponding to superspins $s$ and $s+1/2$, with $s$ a positive integer, and thus they cannot be considered as genuine off-shell ${\cal N}=2$ supermultiplets. The closure of the algebra of ${\cal N}=2$ supersymmetry transformations in these models is a specific feature of the AdS4 supersymmetry, and it is also manifested in the case of ${\cal N}=2$ supersymmetric nonlinear sigma models in AdS4 realized in terms of ${\cal N}=1$ chiral superfields Butter:2011zt ; Butter:2011kf . Genuine off-shell massless ${\cal N}=2$ supermultiplets were constructed in Buchbinder:2021ite within the harmonic superspace approach.
@@
In three dimensions, massless higher-spin supersymmetric models do not have any propagating degrees of freedom. However, these models can be combined with linearized superconformal higher-spin actions [^4] Linearized $\cal N$}-extended superconformal gauge-invariant actions have the following general form Buchbinder:2019yhl : $S^{(n|{\cal N})}[H_{\alpha(n)}]=({\rm i}^{n}/2)\int{\rm d}^{3}x{\rm d}^{2\cal N}\theta\,H^{\alpha(n)}{W}_{\alpha(n)}(H)$, with $n>0$. Here $H_{\alpha(n)}=H_{(\alpha_{1}\dots\alpha_{n})}$ is the unconstrained superconformal gauge prepotential and $W_{\alpha(n)}$ is the linearized super-Cotton tensor, a unique gauge-invariant primary descendant of $H_{\alpha(n)}$. In the non-supersymmetric (${\cal N}=0$) case, $W_{\alpha(n)}$ was constructed in PopeTownsend for even $n$ and Kuzenko:2016qdw for odd $n$. In the ${\cal N}=1$ and ${\cal N}=2$ cases, $W_{\alpha(n)}$ was constructed in Kuzenko:2016qdw ; Kuzenko:2016qwo and KO , respectively. Explicit expressions for $W_{\alpha(n)}$ for ${\cal N}>2$ were given in Buchbinder:2019yhl . to generate off-shell topologically massive higher-spin supersymmetric theories, following the philosophy of topologically massive gauge theories WS ; JS ; DJT1 ; DJT2 ; DeserKay . Recently, there has been much progress in the construction of such off-shell theories with various supersymmetries. It is based on the approaches developed in Kuzenko:2016qwo for ${{\cal N}}=1$ and in KO for ${{\cal N}}=2$ Poincaré supersymmetries. They were later extended to $\rm AdS_{3}$ for the following cases: ${\cal N}=1$ Kuzenko:2018lru ; HK19 ; Kuzenko:2021hyd , ${\cal N}=(1,1)$ HKO and ${\cal N}=(2,0)$ HK18 . Linearized actions for $\cal N$-extended higher-spin superconformal gravity were derived in Buchbinder:2019yhl . [^5] In four dimensions, linearized superconformal higher-spin actions were derived in KMT ; KP ; Kuzenko:2020jie for ${\cal N}=1$ and Kuzenko:2021pqm for ${\cal N}>1$, including the discovery of ${\cal N}=2$ superconformal gravitino multiplet HKR . The ${\cal N}=2$ results of Kuzenko:2021pqm ; HKR were recast in terms of projective Kuzenko:2024vms and harmonic BIZ24 ; Ivanov:2024bsb prepotentials.
@@
The rest of the paper is organized as follows. In Sec. 2 we give a short overview of existing superfield approaches to 5D supergravity and discuss some key results which represent a background for further sections. In Sec. 3, we introduce prepotentials for 5D Poincaré supergravity in the harmonic superspace framework. We use these prepotentials for construction of the superfield action for linearized supergravity in flat Minkowski superspace and consider a minimal hypermultiplet coupling to this supergravity background. This allows us to derive conserved current superfields for the $q$-hypermultiplet model. This section serves as a starting point for the introduction of superfield prepotentials with the corresponding gauge transformations for higher-spin supermultiplets which are given in Sec. 4. There we construct gauge-invariant quadratic actions for higher-spin supermultiplets in harmonic superspace and derive the corresponding superfield equations of motion. Next, we build a cubic interaction vertex for these supermultiplets with the $q$-hypermultiplet. This vertex allows us to derive conserved higher currents in the $q$-hypermultiplet model. Section 5 is devoted to a discussion of other interesting open problems which may be addressed by building upon the results of the present paper. In Appendix A we give a summary of our superspace notation and conventions, while Appendix B contains technical details of the derivation of the component structure of higher-spin prepotentials in the Wess-Zumino gauge. Finally, in Appendix C, we present prepotentials and gauge transformations for 5D conformal supergravity.
@@ 2 Comments on 5D superfield supergravity
General ${\cal N}=1$ supergravity-matter systems in five dimensions are naturally described using the superconformal tensor calculus Ohashi1 ; Ohashi2 ; Ohashi3 ; Ohashi4 ; Bergshoeff1 ; Bergshoeff2 ; Bergshoeff3 (see also Lauria:2020rhc for a pedagogical review [^7] In the literature, ${\cal N}=1$ supersymmetry in five dimensions is often denoted ${\cal N}=2$, see e.g. Lauria:2020rhc ; Ohashi1 ; Ohashi2 ; Ohashi3 ; Ohashi4 ; Bergshoeff1 ; Bergshoeff2 ; Bergshoeff3 , by analogy with the 4D ${\cal N}=2$ case to indicate that the number of supercharges is equal to eight.) in which hypermultiplets are either on-shell or involve a gauged central charge. These hypermultiplet realizations cannot be used to provide an off-shell formulation for the most general locally supersymmetric sigma model. Such a sigma model formulation is known to require one to use off-shell hypermultiplets possessing an infinite number of auxiliary fields, which make them extremely difficult to work with at the component level. This technical problem was solved within the superspace approach to 5D ${\cal N}=1$ supergravity-matter systems KT-M_5D2 ; KT-M_5D3 ; KT-M08 by putting forward the concept of covariant projective multiplets. These supermultiplets are a curved superspace extension of the 4D ${\cal N}=2$ and 5D ${\cal N}=1$ superconformal projective multiplets Kuzenko2006 ; K07 . The latter reduce to the off-shell projective multiplets pioneered by Lindström and Roček KLR ; LR1 ; LR2 in the 4D ${\cal N}=2$ super-Poincaré case and generalized to the cases of 5D ${\cal N}=1$ Poincaré and AdS supersymmetries in KL and Kuzenko:2007aj , respectively.
The powerful virtues of the superconformal tensor calculus and of the superfield supergravity are naturally combined in the framework of 5D ${\cal N}=1$ conformal superspace Butter:2014xxa which was obtained by gauging the 5D superconformal algebra in superspace. This work reproduced practically all off-shell constructions derived at the time and provided a plethora of new results. In particular, it offered covariant superspace realizations for all off-shell formulations of 5D ${\cal N}=1$ supergravity with the Weyl supermultiplet coupled to two compensating supermultiplets. [^8] Numerous applications of the 5D conformal superspace approach have appeared in Refs. Hutomo:2022hdi ; Gold:2023dfe ; Gold:2023ymc ; Gold:2023ykx ; Gold:2025ttt . The supergravity equations of motion were obtained in Ref. Butter:2014xxa by making use of the fact that the 5D ${\cal N}=1$ Weyl multiplet is described in superspace by a single unconstrained real prepotential, $\mathfrak{U}$, similarly to the 4D ${\cal N}=2$ case HST ; KT ; Butter:2010sc . Given a supersymmetric theory of (matter) superfields $\varphi^{i}$ coupled to the Weyl multiplet, the supercurrent of this theory is a (dimension-3 primary) real scalar superfield defined by
@@
In the case of 4D ${\cal N}=2$ supergravity, the unconstrained prepotential $\mathfrak{U}$ for the Weyl multiplet originates most elegantly KT ; Butter:2010sc within the harmonic superspace approach GIKOS ; Galperin:1987em ; Galperin:1987ek . The same approach has recently been used to construct linearized off-shell actions for 4D ${\cal N}=2$ massless higher-spin supermultiplets Buchbinder:2021ite ; Buchbinder:2022kzl ; Buchbinder:2022vra since the supergravity prepotentials in the harmonic superspace possess a natural generalization to higher-spin supermultiplets. Therefore, in the next section, we develop prepotentials for linearized ${\cal N}=1$ Poincaré supergravity in 5D harmonic superspace and construct the corresponding quadratic action for the spin-2 supermultiplet. [^9] In this paper, a supermultiplet which contains bosonic fields of integer spins $s$ and s1s-1, as well as one Dirac-type fermionic field of spin s1/2s-1/2 on shell is referred to as a spin-$s$ supermultiplet. In Section 4, this formulation will be extended for the description of higher-spin multiplets in 5D harmonic superspace. The hypermultiplet coupling to the 5D Weyl multiplets is discussed in Appendix C.
@@
As mentioned above, off-shell formulations for 5D minimal supergravity are obtained by coupling the Weyl multiplet to two compensating supermultiplets. One of the compensators is a vector supermultiplet. In the harmonic superspace framework, prepotentials for the linearized Weyl multiplet are given by the superfields $h^{++m}$, $h^{++\hat{\alpha}+}$ and $h^{(+4)}$ with constraints and gauge transformations described in Appendix C. The coupling to the vector multiplet compensator is accounted for by adding the prepotential $h^{++5}$ such that the covariant harmonic derivative becomes
@@
where $D^{++}$ and $D^{--}$ are flat harmonic derivatives and $\partial_{5}$ is the central charge operator. The functional form of $h^{++5}$ is fixed by the requirement that the operator (2) preserves the analyticity of a massive $q$-hypermultiplet (with a non-vanishing central charge).
@@
which are superconformal provided that $L_{\mathbf{a}}^{i}$ is primary and dimensionless. The action for 5D ${\cal N}=1$ Poincaré supergravity in this setting is proportional to the vector multiplet action with the wrong sign. Several equivalent expressions for the supergravity action exist in the literature, including the following Butter:2014xxa :
@@
where $V_{ij}$ is Mezincescu’s prepotential [^11] Within the conformal superspace framework of Butter:2014xxa , Mezincescu’s prepotential Z2 , $V_{ij}=V_{ji}$, is a real primary superfield of dimension 2-2 defined modulo transformation (6.50) in Butter:2014xxa . for the vector multiplet, $W$ is the field strength of the vector multiplet, and $\nabla^{ij}=\nabla^{\hat{\alpha}(i}\nabla_{\hat{\alpha}}^{j)}$, see Butter:2014xxa for the technical details. The nonlinear multiplet appears in the action (4) implicitly in the sense that the constraints (3) express some component fields of the Weyl multiplet in terms of those belonging to the nonlinear multiplet. The curved superspace geometry, which corresponds to the supergravity formulation under consideration, is completely specified by two curvature tensors: a spinor superfield $\Xi_{\hat{\alpha}}^{i}$ and a super Weyl tensor $W_{ab}=-W_{ba}$. [^12] This geometry may be obtained from 5D conformal superspace by analogy with the 4D ${\cal N}=2$ analysis of Kuzenko:2025bud , and is similar to the 4D ${\cal N}=2$ superspace geometry introduced in Howe ; Castellani:1980cu ; Gates . The supergravity equation of motion corresponding to (4) is
$\Xi_{\hat{\alpha}}^{i}=0~.$ (5)
In the harmonic superspace framework GIOS ; Galperin:1987ek , the non-linear multiplet is described by an analytic superfield $N^{++}$ satisfying the nonlinear constraint [^13] The relationship between $N^{++}$ and $L_{\mathbf{a}}^{i}$ is described in Butter:2015nza .